OpenStudy (anonymous):

hi when u do fractions that have add and minus in themdo u do it as one sum or split it into two e.g 2/5+3/6-2/3 =

OpenStudy (anonymous):

2/5+3/6-2/3 = 12+15-20/30 = 7/30 note that 30 is the common factor on 5, 6 and 3... so u divide 30 by each denominator then multiply result by numerator n write that number...

OpenStudy (anonymous):

thanks for the explanation itsmuch clearer now

OpenStudy (anonymous):

you are welcome